OpenStudy (anonymous):

Could someone check my answer and tell me if its right? I got 50.7 V for this question The peak value of an AC passing through a 13.0 W device is 3.9 A. What is the rms voltage across it?

OpenStudy (primeralph):

P = I*V, so V = P/I.

OpenStudy (anonymous):

omg i was thinking of V= I * R idk y

OpenStudy (anonymous):

so 3.33